# Compression and Distillation of Data Quadruplets in Non-intrusive Reduced-order Modeling

### v3

@@ 4.2 Using Available Impulse Response Data
In many applications, obtaining frequency-domain measurements is impractical. Instead, impulse response data is frequently utilized for various analysis and design tasks. When direct impulse response measurements are not feasible, a step input can be applied, and the impulse response can be obtained through differentiation. While a detailed review of these methods falls outside the scope of this paper, we refer readers to [32, 33, 34, 35, 36] for further insights. In this subsection, we demonstrate that existing impulse response data is sufficient for non-intrusive data-driven implementation of IRKA.
@@
It is now evident that IRKA can be implemented using impulse response data, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “time-domain quadrature-based IRKA (TD-Quad-IRKA)”, is provided in Algorithm 3.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
7:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
Range of Impulse Response Sampling: Let us restrict the integral range of (45) from $[0,\infty]$ to $[0,t_{f}]$ rad/sec. Then $V_{\tau}=V\big{|}_{0}^{t_{f}}$ solves the following Sylvester equation:
@@
as described in [37]. Theoretically, $e^{E^{-1}At_{f}}\rightarrow 0$ and $e^{-S_{b}t_{f}}\rightarrow 0$ as $t_{f}\rightarrow\infty$. In practice, $e^{E^{-1}At_{f}}$ and $e^{-S_{b}t_{f}}$ rapidly approach zero for a finite $t_{f}$, depending on how far the eigenvalues of $E^{-1}A$ and $-S_{b}$ are from the $j\omega$-axis. The farther the eigenvalues of $E^{-1}A$ and $-S_{b}$ are from the $j\omega$-axis, the faster the exponentials $e^{E^{-1}At_{f}}$ and $e^{-S_{b}t_{f}}$ decay to zero. As a result, $V_{\tau}$ becomes numerically equivalent to $V$ beyond a finite time range. Therefore, in practice, the nodes of the numerical quadrature can be confined to a finite time range, especially when the poles of $G(s)$ are located far from the $j\omega$-axis in the left half of the $s$-plane. Consequently, we can use a finite $t_{f}$ in the numerical quadrature rule, and the integration limits can be mapped accordingly. For instance, the integration limits $[-1,1]$ in the Gauss-Legendre quadrature rule can be mapped to $[0,t_{f}]$ using the following transformation:
@@ 4.3 Using Available Transfer Function Samples
In this subsection, we illustrate how the block version of PORK can be utilized to develop a non-intrusive, data-driven implementation of IRKA using available transfer function samples. Recall the following expressions:
[…]
Similar to the ADI method, if we substitute $G(s)$ in (52) with its interpolant generated by the block version of I-PORK at the available interpolation points $(\alpha_{1},\cdots,\alpha_{n_{p}})$, and replace $G(s)$ in (53) with its interpolant produced by the block version of O-PORK at the interpolation points $(\beta_{1},\cdots,\beta_{n_{q}})$, we can achieve a non-intrusive, data-driven implementation of IRKA. The interpolation points $\alpha_{i}$ and $\beta_{i}$ are all located in the right-half of the $s$-plane.
@@
[…]
@@
It can then be observed that $V\approx\tilde{V}\hat{V}_{r}$ and $W\approx\tilde{W}\hat{W}_{r}$. Assuming this approximation is exact, the ROM satisfying the interpolation condition (3) can be obtained by reducing the Loewner quadruplet $(E_{\alpha,\beta},A_{\alpha,\beta},B_{\alpha,\beta},C_{\alpha,\beta})=(\tilde{W}^{*}E\tilde{V},\tilde{W}^{*}A\tilde{V},\tilde{W}^{*}B,C\tilde{V})$ as follows:
@@
It is now evident that IRKA can be implemented using available transfer function samples $G(\alpha_{i})$ and $G(\beta_{i})$, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “PORK-based IRKA (PORK-IRKA)”, is provided in Algorithm 4.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
6:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
4.4 Tracking the Error $||G(s)-G_{r}(s)||_{\mathcal{H}_{2}}$
@@
$=$ $||G(s)||_{\mathcal{H}_{2}}^{2}+||G_{r}(s)^{(i-1)}||_{\mathcal{H}_{2}}^{2}-2\text{trace}\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}.$
Thus, with a delay of one iteration, the error $||G(s)-G_{r}(s)||_{\mathcal{H}_{2}}$ can be tracked if $||G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ is computed in every iteration. It is important to note that the original expression presented in [38] is intrusive, whereas the expression above is its non-intrusive equivalent. To summarize, the error in data-driven IRKA can also be monitored non-intrusively by tracking the following term:
$||G_{r}(s)^{(i-1)}||_{\mathcal{H}_{2}}^{2}-2\text{trace}\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}.$
However, it should be noted that the term $-2\,trace\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}$ is an approximation and not exact. Its accuracy depends on the precision of the approximations of the integrals (38) and (39) or (45) and (46).
5 Data-driven Implementations of IRKA for Discrete-time Systems
@@
The necessary conditions for a local optimum of $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ are given by:
@@
Similar to the continuous-time case, since the ROM $G_{r}(z)$ is initially unknown, the discrete-time IRKA (DT-IRKA) [39] uses fixed-point iterations starting from an arbitrary initial guess of the interpolation data to search for a local optimum. After each iteration, the interpolation data is updated as $\sigma_{i}=\mu_{i}=\frac{1}{\hat{\lambda}_{i}}$, $b_{i}=\hat{r}_{i}$, and $c_{i}=\hat{l}_{i}^{*}$ until convergence is achieved. Upon convergence, a local optimum of $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ is achieved.
However, since DT-IRKA updates the interpolation points during the process, it requires evaluating the transfer function samples at these updated points. This necessitates halting DT-IRKA and conducting new experiments to obtain new samples, which is often impractical. Additionally, since $G_{r}(z)$ is stable, the interpolation points $\frac{1}{\hat{\lambda}_{i}}$ lie outside the unit circle. Exciting the system at these frequencies can be dangerous or even impossible [40]. These challenges motivate the development of offline transfer function sampling strategies that utilize existing data.
@@ 5.1 Using Available Frequency Response Data
From these definitions, it is clear that the summations (75) and (76) can be represented as $\tilde{V}\hat{V}_{r}$ and $\hat{W}_{r}^{*}\tilde{W}^{*}$, respectively. Thus, $V\approx\tilde{V}\hat{V}_{r}$ and $W\approx\tilde{W}\hat{W}_{r}$. Let us assume, for a moment, that this approximation is exact. In this case, the ROM satisfying the interpolation condition (3) can be obtained by reducing the Loewner quadruplet $(E_{jw},A_{jw},B_{jw},C_{jw})=(\tilde{W}^{*}E\tilde{V},\tilde{W}^{*}A\tilde{V},\tilde{W}^{*}B,C\tilde{V})$ as follows:
@@
It is now clear that DT-IRKA can be implemented using frequency-domain data $G(e^{j\xi_{i}})$, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. Furthermore, since the eigenvalues $\frac{1}{\sigma_{i}}$ of $\bar{S}_{b}$ lie within the unit circle, the interpolation points can be outside the unit circle, and $G(s)$ and $G^{\prime}(s)$ can be sampled outside the unit circle without any issues. The pseudo-code for the frequency-domain quadrature-based DT-IRKA (FD-Quad-DTIRKA) is outlined in Algorithm 5.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
7:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@ 5.2 Using Available Impulse Response Data
From these definitions, it is evident that the sums (85) and (86) can be represented as $\tilde{V}\hat{V}_{r}$ and $\hat{W}_{r}^{*}\tilde{W}^{*}$, respectively. Thus, we have the approximations $V\approx\tilde{V}\hat{V}_{r}$ and $W\approx\tilde{W}\hat{W}_{r}$. Let us assume, for a moment, that this approximation is exact. In this case, the ROM satisfying the interpolation condition (3) can be obtained by reducing the impulse data quadruplet $(E_{k},A_{k},B_{k},C_{k})=(\tilde{W}^{*}E\tilde{V},\tilde{W}^{*}A\tilde{V},\tilde{W}^{*}B,C\tilde{V})$ as follows:
@@
It is now clear that DT-IRKA can be implemented using impulse response data $h(k)$, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. Additionally, since the eigenvalues $\frac{1}{\sigma_{i}}$ of $\bar{S}_{b}$ lie within the unit circle, the interpolation points can be outside the unit circle, allowing $G(s)$ and $G^{\prime}(s)$ to be sampled outside the unit circle without any issues. The pseudo-code for the time-domain DT-IRKA (TD-DTIRKA) is provided in Algorithm 6.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
7:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@ 6 Pseudo-optimal Rational Krylov (PORK) Algorithm for Discrete-time Systems
In this section, we extend PORK to discrete-time systems and show that the discrete-time version maintains properties comparable to its continuous-time counterpart. Building on the findings from this section, we will formulate non-intrusive, data-driven implementations of BT and DT-IRKA in the following section.
@@ 6.3 Approximation of Gramians
$||G(z)-G_{r-1}(z)||_{\mathcal{H}_{2}}^{2}$ $=||G(z)||_{\mathcal{H}_{2}}^{2}-||G_{r-1}(z)||_{\mathcal{H}_{2}}^{2},$
$||G_{r}(z)-G_{r-1}(z)||_{\mathcal{H}_{2}}^{2}$ $=||G_{r}(z)||_{\mathcal{H}_{2}}^{2}-||G_{r-1}(z)||_{\mathcal{H}_{2}}^{2},$
$||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ $=||G(z)||_{\mathcal{H}_{2}}^{2}-||G_{r}(z)||_{\mathcal{H}_{2}}^{2},$
$||G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ $\geq||G_{r-1}(z)||_{\mathcal{H}_{2}}^{2},$
$||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ $\leq||G(z)-G_{r-1}(z)||_{\mathcal{H}_{2}}^{2}.$
Therefore, as the order of the ROM increases, $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}$ decays monotonically. A similar result can be shown for O-PORK.
@@ 7 Non-intrusive PORK-based Low-rank Balanced Truncation for Discrete Time Systems
The quality of approximation of $P$ and $Q$ can be tracked non-intrusively by observing the growth of $CV\bar{Q}_{s}^{-1}V^{*}C^{T}$ and $B^{T}W^{*}\bar{P}_{s}^{-1}W^{*}B$, respectively. Note that $CV$, $W^{*}B$, $P_{r}=\bar{Q}_{s}^{-1}$, and $Q_{r}=\bar{P}_{s}^{-1}$ can be computed using interpolation data and samples of $G(z)$ at the interpolation points $\sigma_{i}$ and $\mu_{i}$. Furthermore, since $W^{*}EV$ and $W^{*}AV$ can also be computed non-intrusively from (27) via the Loewner framework, a data-driven low-rank BT algorithm can be formulated, analogous to its continuous-time counterpart. The pseudo-code for the data-driven PORK-based discrete-time BT (DD-PORK-DTBT) is presented in Algorithm 7.
Algorithm 7 DD-PORK-DTBT

### v4

@@ 4.2 Using Available Impulse Response Data
In many applications, obtaining frequency-domain measurements is impractical. Instead, impulse response data is frequently utilized for various analysis and design tasks. When direct impulse response measurements are not feasible, a step input can be applied, and the impulse response can be obtained through differentiation. While a detailed review of these methods falls outside the scope of this paper, we refer readers to [30, 31, 32, 33, 34] for further insights. In this subsection, we demonstrate that existing impulse response data is sufficient for non-intrusive implementation of IRKA.
@@
It is now evident that IRKA can be implemented using impulse response data, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “time-domain quadrature-based IRKA (TD-Quad-IRKA)”, is provided in Algorithm 3.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
7: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
Range of Impulse Response Sampling: Let us restrict the integral range of (46) from $0,\infty]$ to $0,t_{f}]$ rad/sec. Then $V_{\tau}=V\big|_{0}^{t_{f}}$ solves the following Sylvester equation:
@@
as described in [35]. Theoretically, $e^{E^{-1}At_{f}}\rightarrow 0$ and $e^{-S_{b}t_{f}}\rightarrow 0$ as $t_{f}\rightarrow\infty$. In practice, $e^{E^{-1}At_{f}}$ and $e^{-S_{b}t_{f}}$ rapidly approach zero for a finite $t_{f}$, depending on how far the eigenvalues of $E^{-1}A$ and $S_{b}$ are from the $j\omega$-axis. The farther the eigenvalues of $E^{-1}A$ and $S_{b}$ are from the $j\omega$-axis, the faster the exponentials $e^{E^{-1}At_{f}}$ and $e^{-S_{b}t_{f}}$ decay to zero. As a result, $V_{\tau}$ becomes numerically equivalent to $V$ beyond a finite time range. Therefore, in practice, the nodes of the numerical quadrature can be confined to a finite time range, especially when the poles of $G(s)$ are located far from the $j\omega$-axis in the left half of the $s$-plane. Consequently, we can use a finite $t_{f}$ in the numerical quadrature rule, and the integration limits can be mapped accordingly. For instance, the integration limits [1,1][-1,1] in the Gauss-Legendre quadrature rule can be mapped to $0,t_{f}]$ using the following transformation:
@@
Avoiding the Samples of $h^{\prime}(t)$: Similar to the frequency domain, by setting $S_{b}$ and $L_{b}$ as in (45), we can compute the following:
$\begin{bmatrix}G(\sigma_{1})&G^{\prime}(\sigma_{1})&\cdots&G(\sigma_{r})&G^{\prime}(\sigma_{r})\end{bmatrix}\approx\begin{bmatrix}h(t_{1})&\cdots&h(t_{n_{p}})\end{bmatrix}\hat{V}_{r}.$
Again, this results in some loss of accuracy, as single-sided moment matching is less accurate than double-sided moment matching.
@@ 4.3 Using Available Transfer Function Samples
In this subsection, we illustrate how the block version of PORK can be utilized to develop a non-intrusive implementation of IRKA using available transfer function samples. Recall the following expressions:
$CV$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}G(j\nu)L_{b}(-j\nu I+S_{b})d\nu,$ (53)
$W^{*}B$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\nu I+S_{c})^{-1}L_{c}G(j\nu)d\nu.$ (54)
Similar to the ADI method, if we substitute $G(s)$ in (53) with its interpolant generated by the block version of I-PORK at the available interpolation points $(\alpha_{1},\cdots,\alpha_{n_{p}})$, and replace $G(s)$ in (54) with its interpolant produced by the block version of O-PORK at the interpolation points $(\beta_{1},\cdots,\beta_{n_{q}})$, we can achieve a non-intrusive implementation of IRKA. The interpolation points $\alpha_{i}$ and $\beta_{i}$ are all located in the right-half of the $s$-plane.
@@
$CV$ $\approx\frac{1}{2\pi}C\tilde{V}\Big(\int_{-\infty}^{\infty}(j\nu I-A_{\alpha})^{-1}B_{\alpha}L_{b}(-j\nu I+S_{b})d\nu\Big),$ (62)
$W^{*}B$ $\approx\Big(\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\nu I+S_{c})^{-1}L_{c}C_{\beta}(j\nu I-A_{\beta})^{-1}d\nu\Big)\tilde{W}^{*}B.$ (63)
@@
[…]
@@
It is now evident that IRKA can be implemented using available transfer function samples $G(\alpha_{i})$ and $G(\beta_{i})$, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “PORK-based IRKA (PORK-IRKA)”, is provided in Algorithm 4.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
6: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
4.4 Tracking the Error $G(s)-G_{r}(s)||_{\mathcal{H}_{2}}$
@@
[…]
Thus, with a delay of one iteration, the error $G(s)-G_{r}(s)||_{\mathcal{H}_{2}}$ can be tracked if $G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ is computed in every iteration. It is important to note that the original expression presented in [36] is intrusive, whereas the expression above is its non-intrusive equivalent. To summarize, the error in non-intrusive IRKA can also be monitored non-intrusively by tracking the following term:
[…]
However, it should be noted that the term $2\,trace\Big(C_{r}^{(i)}\big(C_{r}^{(i-1)}T_{r}^{(i-1)}\big)^{*}\Big)$ is an approximation and not exact. Its accuracy depends on the precision of the approximations of the integrals (38) and (39) or (46) and (47).
5 Non-intrusive Implementations of IRKA for Discrete-time Systems
@@
The necessary conditions for a local optimum of $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ are given by:
@@
Similar to the continuous-time case, since the ROM $G_{r}(z)$ is initially unknown, the discrete-time IRKA (DT-IRKA) [37] uses fixed-point iterations starting from an arbitrary initial guess of the interpolation data to search for a local optimum. After each iteration, the interpolation data is updated as $\sigma_{i}=\mu_{i}=\frac{1}{\hat{\lambda}_{i}}$, $b_{i}=\hat{r}_{i}$, and $c_{i}=\hat{l}_{i}^{*}$ until convergence is achieved. Upon convergence, a local optimum of $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ is achieved.
However, since DT-IRKA updates the interpolation points during the process, it requires estimating the transfer function samples at these updated points. This necessitates halting DT-IRKA and conducting new experiments to estimate new samples, which is often impractical.
@@ 5.1 Using Available Frequency Response Data
[…]
@@
It is now clear that DT-IRKA can be implemented using frequency-domain data $G(e^{j\xi_{i}})$, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. The pseudo-code for the frequency-domain quadrature-based DT-IRKA (FD-Quad-DTIRKA) is outlined in Algorithm 5.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
7: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@ 5.2 Using Available Impulse Response Data
[…]
@@
It is now clear that DT-IRKA can be implemented using impulse response data $h(k)$, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. The pseudo-code for the time-domain DT-IRKA (TD-DTIRKA) is provided in Algorithm 6.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
7: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@ 6 Pseudo-optimal Rational Krylov (PORK) Algorithm for Discrete-time Systems
In this section, we extend PORK to discrete-time systems and show that the discrete-time version maintains properties comparable to its continuous-time counterpart. Building on the findings from this section, we will formulate non-intrusive implementations of BT and DT-IRKA in the following section.
@@ 6.3 Approximation of Gramians
[…]
Therefore, as the order of the ROM increases, $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}$ decays monotonically. A similar result can be shown for O-PORK.
@@ 7 Non-intrusive PORK-based Low-rank Balanced Truncation for Discrete Time Systems
The quality of approximation of $P$ and $Q$ can be tracked non-intrusively by observing the growth of $CV\bar{Q}_{s}^{-1}V^{*}C^{T}$ and $B^{T}W^{*}\bar{P}_{s}^{-1}W^{*}B$, respectively. Note that $CV$, $W^{*}B$, $P_{r}=\bar{Q}_{s}^{-1}$, and $Q_{r}=\bar{P}_{s}^{-1}$ can be computed using interpolation data and samples of $G(z)$ at the interpolation points $\sigma_{i}$ and $\mu_{i}$. Furthermore, since $W^{*}EV$ and $W^{*}AV$ can also be computed non-intrusively from (27) via the Loewner framework, a non-intrusive low-rank BT algorithm can be formulated, analogous to its continuous-time counterpart. The pseudo-code for the non-intrusive PORK-based discrete-time BT (NI-PORK-DTBT) is presented in Algorithm 7.
Algorithm 7 NI-PORK-DTBT